What is 2 over 43 equal j over 13. How do you solve that and do u have to reduce?: What is 2 over 43 equal j over 13. How do you solve that and do u have to reduce?
Answer by nabla(409)

2/43=j/13
Multiply both sides by 13.
26/43=j
This is reduced.